OpenStudy (loser66):

Someone helps me please, how to calculate:

OpenStudy (loser66):

They give me \(E_{10} = 50521\) but no matter how hard I tried, I couldn't get that answer.

OpenStudy (loser66):

\[E_{2n}-\left(\begin{matrix}2n\\2n-2\end{matrix}\right)E_{2n-2}+\left(\begin{matrix}2n\\2n-4\end{matrix}\right)E_{2n-4}+\cdots+(-1)^n\left(\begin{matrix}2n\\2\end{matrix}\right)E_{2}+(-1)^n =0\]

OpenStudy (loser66):

\[E_{2n}=\left(\begin{matrix}2n\\2n-2\end{matrix}\right)E_{2n-2}-\left(\begin{matrix}2n\\2n-4\end{matrix}\right)E_{2n-4}-\cdots-(-1)^n\left(\begin{matrix}2n\\2\end{matrix}\right)E_{2}-(-1)^n \]

OpenStudy (loser66):

\(E_0 =1\)

OpenStudy (loser66):

If n =1 \[E_2= -(-1)^1 =1\]

OpenStudy (loser66):

\(E_4= (-1)^2\left(\begin{matrix}4\\2\end{matrix}\right)E_2-(-1)^2\\E_4=6E_2-1 = 5\)

OpenStudy (loser66):

if n=3 \[E_6=\left(\begin{matrix}6\\4\end{matrix}\right)E_4-(-1)^3\left(\begin{matrix}2n\\2\end{matrix}\right)E_{2}-(-1)^3\\E_6 = 15E_4+15E_2+1=91 \]

OpenStudy (loser66):

But they said E6 =61, how?